# A mathematical framework of intelligence and consciousness based on Riemannian Geometry

### v3

@@ 2 Thought flow and Geodesic
Concept: Concept: In cognitive science, we propose that without perturbation or external stimulation, the thought flow navigates naturally on curved manifolds, following the geometry of the space. This unperturbed path satisfies the definition of a geodesic. In this context, the geodesic path is represented as a sequence of tokens activated along this path, reflecting the natural trajectory of thought flow in the high-dimensional token space. The perturbed thought flow that is ”forced” by stimuli or external input to deviate from its geodesic, and is the most common scenario, will be analysed in the later section of the Consciousness. The geodesic is the path that minimises the distance between points, analogous to the shortest path on a curved surface. In this theory, the geodesic $\gamma(t)$ represents the state of intelligence at time $t$.
@@ 4 Attention Mechanism
The curvature and structure of the geodesic are results of the training process, capturing the intrinsic links, organisation, and distribution of tokens. The attention mechanism measures the contextual significance and tokens’ correlation, which is geometrically represented by the manifold’s curvature and structure. This creates a fundamental link between the attention mechanism and the tangent vector. In mathematical terms, while the tangent vector provides a local derivative (instantaneous change), the attention mechanism offers a global perspective by integrating the influence of prior tokens. This integrated influence aligns with the geodesic’s learned curvatures, as both are products of the underlying manifold’s geometry.
@@ 5 Context Vector and Contextual Embedding
$h_{t,j}=c_{i}=\sum_{j}\alpha_{ij}(W\phi v^{\prime}_{j})$
Here, $c_{i}$ is the context vector for token $t_{i}$, $\alpha_{ij}$ are the attention weights, and $W\phi$ is the learned value matrix.
@@ 6 Predicted Token
The predicted token concept is derived from the context vector, representing the next token in the thought flow influenced by contextual information. This concept can be expressed mathematically by two different equations:
@@ 8 Geodesic Equation with Feedback
  • $\kappa$ is the consciousness intensity index of the predicted error (feedback), applied as a modulator to the prediction error. It represents the intensity with which the agent processes the prediction error. A higher $\kappa$ indicates a more intense response to the prediction error.
@@
This describes the natural geodesic path of the thought flow without any correction from prediction error feedback. The system follows its expected trajectory, indicating a stable and accurate state of consciousness.
@@ 9.2 Consciousness Threshold
The current theory of the geometry of intelligence is grounded in the concept of tokens as fundamental elements of the intelligence space. These tokens form manifolds with geometric properties such as curvature, and their organisation and activation sequences provide a comprehensive structure and dynamics of intelligence, respectively. In this framework, tokens represent discrete units of information that can be multi-module tokens (Radford et al., 2021; Lyu et al., 2023) embedded in a high-dimensional space. These tokens are not only organised by their inherent features but also by their roles and relationships in various cognitive processes.
The geometric representation of these tokens forms manifolds that capture both the static distribution of features and the dynamic sequences of activation. The curvature of these manifolds reflects the complexity and interconnections within the cognitive processes. The static aspect of the theory encompasses the organisation of tokens based on their features, similar to the learned manifolds in VAEs and GANs, providing a foundational structure that captures the intrinsic relationships between different elements of intelligence.
@@
Feedback mechanisms in this theory adjust the manifold’s structure based on non-zero prediction errors and external inputs, facilitating learning and adaptation. In Claude 3 Sonnet, feature manipulation and subsequent response changes reflect a feedback mechanism where the model adjusts its internal representations.
@@
    $\frac{d^{2}\gamma^{\mu}(t)}{dt^{2}}+\Gamma^{\mu}_{\nu\lambda}\frac{d\gamma^{\nu}(t)}{dt}\frac{d\gamma^{\lambda}(t)}{dt}=\kappa\frac{d^{2}\psi(\Delta^{\mu}(t))}{dt^{2}}+\eta\frac{d^{2}I^{\mu}(t)}{dt^{2}}$
@@
$\frac{d^{2}\gamma^{\mu}(t)}{dt^{2}}+\Gamma^{\mu}_{\nu\lambda}\frac{d\gamma^{\nu}(t)}{dt}\frac{d\gamma^{\lambda}(t)}{dt}=\kappa\cdot\frac{d^{2}\psi(\Delta^{\mu}(t))}{dt^{2}}+\eta\cdot\frac{d^{2}I^{\mu}(t)}{dt^{2}}$

### v4

@@ 2 Thought flow and Geodesic
Concept: In cognitive science, we propose that without perturbation or external stimulation, the thought flow navigates naturally on curved manifolds, following the geometry of the space. This unperturbed path satisfies the definition of a geodesic. In this context, the geodesic path is represented as a sequence of tokens activated along this path, reflecting the natural trajectory of thought flow in the high-dimensional token space. The perturbed thought flow that is ”forced” by stimuli or external input to deviate from its geodesic, and is the most common scenario, will be analysed in the later section of the Consciousness. The geodesic is the path that minimises the distance between points, analogous to the shortest path on a curved surface. In this theory, the geodesic $\gamma(t)$ represents the state of intelligence at time $t$.
@@ 4 Attention Mechanism
The curvature and structure of the manifolds are results of the training process, capturing the intrinsic links, organisation, and distribution of tokens. The attention mechanism measures the contextual significance and tokens’ correlation, which is geometrically represented by the manifold’s curvature and structure. This creates a fundamental link between the attention mechanism and the tangent vector. In mathematical terms, while the tangent vector provides a local derivative (instantaneous change), the attention mechanism offers a global perspective by integrating the influence of prior tokens. This integrated influence aligns with the geodesic’s learned curvatures, as both are products of the underlying manifold’s geometry.
@@ 5 Context Vector and Contextual Embedding
$h_{t,j}=c_{i}=\sum_{j}\alpha_{ij}(W\omega v^{\prime}_{j})$
Here, $c_{i}$ is the context vector for token $t_{i}$, $\alpha_{ij}$ are the attention weights, and $W\omega$ is the learned value matrix.
@@ 6 Predicted Token
This concept can be expressed mathematically by two different equations:
@@ 8 Geodesic Equation with Feedback
  • $\kappa$ is the consciousness intensity index applied as a modulator to the prediction error. It represents the intensity with which the agent processes the prediction error. A higher $\kappa$ indicates a more intense response to the prediction error.
@@
This describes the natural geodesic path of the thought flow without any correction from prediction error feedback. The system follows its expected trajectory.
@@ 9.2 Consciousness Threshold
The tokens (Radford et al., 2021; Lyu et al., 2023) with geometric representation forms manifolds that capture both the static distribution of features and the dynamic sequences of activation. The curvature of these manifolds reflects the complexity and interconnections within the cognitive processes. The static aspect of the theory encompasses the organisation of tokens based on their features, similar to the learned manifolds in VAEs and GANs, providing a foundational structure that captures the intrinsic relationships between different elements of intelligence.
@@
Feedback mechanisms in this theory adjust the manifold’s structure based on non-zero prediction errors and external inputs, facilitating learning and adaptation. In Claude 3 Sonnet, feature manipulation by human/external input acts as a force to adjust the model’s internal representations.
@@
    $\frac{d^{2}\gamma^{\mu}(t)}{dt^{2}}+\Gamma^{\mu}_{\nu\lambda}\frac{d\gamma^{\nu}(t)}{dt}\frac{d\gamma^{\lambda}(t)}{dt}=\kappa\frac{d^{2}\psi(\Delta^{\mu}(t))}{dt^{2}}$